OpenStudy (anonymous):

What is the standard form equation of the line passing through (1, –5) with a slope of –1

jimthompson5910 (jim_thompson5910):

y = mx + b y = -1x + b ... plug in m = -1 -5 = (-1)(1) + b ... plug in x = 1 and y = -5 -5 = -1 + b Now solve for b. Tell me what you get.

OpenStudy (anonymous):

4?? I'm no good in this :/

jimthompson5910 (jim_thompson5910):

close

jimthompson5910 (jim_thompson5910):

y = mx + b y = -1x + b ... plug in m = -1 -5 = (-1)(1) + b ... plug in x = 1 and y = -5 -5 = -1 + b -5+1 = -1 + b +1 -4 = b b = -4

jimthompson5910 (jim_thompson5910):

keep in mind that -5 + 1 = -4 and not 4

jimthompson5910 (jim_thompson5910):

so you were close

OpenStudy (anonymous):

okay I just redid it and got that thank you do much!

jimthompson5910 (jim_thompson5910):

ok, so we go from y = mx+b to y = -x-4

jimthompson5910 (jim_thompson5910):

To get into standard form, we just move that x term to the left side

jimthompson5910 (jim_thompson5910):

y = -x-4 y+x = -x-4+x x+y = -4

jimthompson5910 (jim_thompson5910):

So the answer is x+y = -4

OpenStudy (anonymous):

thnx so very much for your help!

jimthompson5910 (jim_thompson5910):

you're welcome